How many cubical blocks, each with edges of length 4 centimeters, are needed to fill a rectangular box that has inside dimensions 20 centimeters by 24 centimeters by 32 centimeters?
Answer by mananth(16946)

rectangular box inside dimensions 20 centimeters by 24 centimeters by 32
centimeters
Volume V = 20*24*32= 15360 cm^3
volume of cube A = 4^3
A= 64 cm^3
Number of cubes = volume of rectangular box/volume of cube
=15630/64
= 244 cubes